Convert 2.785x10^5 in to km

Convert 2.785x10^5 in to km
Answer #1

Solution:

1 inch = 2.54 * 10^-5 km

2.785 * 10^5 in = 2.785 * 10^5 * 2.54 * 10^-5 km

=> 7.7039 km

thanks.